# The variational principle for a marked Gibbs point process with infinite-range multibody interactions

### v1

@@ 1. Introduction
The Gibbs variational principle is a corner stone in the equilibrium theory of spatially interacting particle systems. In a nutshell, according to the laws of thermodynamics, equilibrium states of a system (subject to thermodynamic principles) must minimize the free energy, which reflects the competition between minimizing energy and maximizing entropy.
@@
On the other hand, equilibrium systems of interacting particles are often described by their conditional finite-volume distributions, in terms of Boltzmann weights with respect to some boundary conditions. Under a consistency condition on these finite-volume measures, one can then hope to find infinite-volume measures that are still consistent with the finite-volume description, and are able to properly describe the macroscopic behavior of the system. This infinite-volume consistency is described via the so-called DLR equations, and ensuring the existence of such measures is one the key tasks in statistical mechanics. In particular, the DLR equations can potentially be solved by multiple infinite-volume measures, constituting a phase transition of non-uniqueness of equilibrium states. However, for a measure to solve the DLR equations and also minimize the free energy is not automatic.
@@
In the case of lattice systems, under the assumption of translation invariance, this link between the DLR equations and the thermodynamic description of equilibria has been established under quite general assumptions for a very large class of models, see [Geo11, Pre76, LR69]. However, moving towards continuum systems, where particles are placed in Euclidean space and even particle numbers in fixed volumes are allowed to fluctuate, the literature becomes much more sparse and usually only deals with particular case studies. In [Geo94, Geo95] the Gibbs variational principle is derived for bounded and unbounded pair potentials in the context of large-deviation results, and with the goal to establish equivalence of ensembles. In particular, no higher-order interactions are allowed, and the particles may not carry individual marks that would for example allow to model interactions between particles of varying sizes. Next, in  [DG09] the Gibbs variational principle is established for a particular class of planar models in which (unmarked) particles interact via their associated Delaunay triangulation. More recently,  [Der16] considers the case of Gibbs point processes with a finite-range interaction and where the existence of the energy density is guaranteed. However, establishing the existence of the energy density is not always easy and needs to be shown specifically for the model at hand. Moreover, many interesting models do not have a finite-range interaction.
@@
The contribution of the present manuscript is therefore to provide a first example of a Gibbs variational principle in the continuum for a system of particles with unbounded marks and with a multibody interaction in which also an unbounded number of particles may jointly interact.
@@
We focus on the Asakura–Oosawa model with depletion interaction [AO54, AO58], which may be informally described as follows. Imagine a volume with a random number of balls of a random size. The balls cannot penetrate each other, leading to a hardcore repulsion. Additionally, there are second-class particles with a fixed size which may penetrate each other without restriction, but not the balls of the first type. Integrating out the Poisson point process of second-class particles, it can be observed that the balls of type one must additionally obey an area interaction: for them to take up more space comes with the additional cost that, in that space, no second-class particles are present. For more details on the relevance of this colloid-polymer mixture model, see [Vri76, BVS14, RED00, LTV24]. Let us mention that this model can be seen as a generalization of the Widom–Rowlinson model [WR70, CCK95, DH19, DH01, JK17] with an additional internal hardcore. The existence of infinite-volume Gibbs measures (in the sense of the DLR equations) for the Asakura–Oosawa model can be deduced from the criteria presented in [RZ20]. Recently, the model has been analyzed more extensively, for example via cluster-expansion techniques [JT20a, JT20b, JLW22] and with respect to the corresponding Langevin dynamics [FKRZ24].
@@ 2.1. Point process formalism
$N_{\Lambda}^{f}:=N_{\Lambda,[0,\infty)}^{f},\quad N_{\Lambda,A}:=N_{\Lambda,A}^{1}\quad\text{ and }\quad N_{\Lambda}:=N_{\Lambda}^{1}.$
@@
where $\mathfrak{B}_{b}$ denotes the set of bounded Borel sets. We also fix some special sets of the position space $\mathbb{R}^{d}$ which will be denoted by
$\Lambda_{n}:=\left[-n/2,n/2\right)^{d},n\in\mathbb{N},\quad\text{and }\qquad C:=\Lambda_{1}$
@@ 2.2. Colloidal multi-body interaction
In other words, the interaction models colloidal particles of radius $R_{i}$ centred at $x_{i}$ in a bath of polymers of size $r$. Note that, by inclusion-exclusion, the area-interaction term can be rewritten in terms of $k$-body potentials as
$H^{\rm ar}(\omega)=\sum_{k\geq 1}\phi_{k}(\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{k}),$
@@ 2.3. Gibbsian formalism
Based on the energy, we now define the finite-volume equilibrium Gibbs measures with respect to suitable boundary conditions. For this, let $\pi^{z}_{\Lambda}$ denote the iid-marked Poisson point process on $\Lambda\times[0,\infty)$ with intensity measure $z\mathds{1}\{x\in\Lambda\}\mathrm{d}x\otimes\mathcal{R}(\mathrm{d}r)$, with $\Lambda\in\mathfrak{B}(\mathbb{R}^{d})$ and $\mathcal{R}$ a probability measure on $[0,\infty)$. We write $\pi^{z}:=\pi^{z}_{\mathbb{R}^{d}}$. Throughout the paper, we make the following integrability assumption,
$\int\mathrm{e}^{r^{d+\delta}}\mathcal{R}(\mathrm{d}r)<\infty\qquad\text{ for some }\delta>0.$
@@
Let us note that, since $H_{\Lambda,\zeta}\geq 0$ and $H_{\Lambda,\zeta}(\emptyset)=0$, we have that $\exp(-z|\Lambda|)\leq Z_{\Lambda,z,\beta,\zeta}\leq 1$ and thus the finite-volume Gibbs measures are well-defined for all boundary conditions.
@@
A probability measure $P\in\mathcal{P}_{\Theta}$ is called an infinite-volume Gibbs measure for $z,\beta>0$ if the following DLR equation holds for any bounded $\Lambda\in\mathfrak{B}(\mathbb{R}^{d})$ and any measurable function $f\geq 0$,
@@
We write $\mathcal{G}_{z,\beta}(H)$ for the set of all infinite-volume Gibbs measures for $z,\beta>0$.
Let us already note, that our main result will in particular establish existence of infinite-volume Gibbs measures for the model defined in 2.1. However, the existence could also be verified by checking the conditions presented in [RZ20].
@@ 2.6. The pressure
$\psi(z,\beta):=\lim_{n\to\infty}\frac{\log(\pi^{z}_{\Lambda_{n}}\left[\mathrm{e}^{-\beta H_{\Lambda_{n},\zeta}}\right])}{|\Lambda_{n}|}.$
Here, the existence of the limit is also quite subtle and depends on the choice of the boundary condition $\zeta$. We will show that, for a class of sufficiently nice boundary conditions, it exists and actually does not depend on the boundary condition itself.
@@ 3. Main result
    A measure $P\in\mathcal{P}_{\Theta}$ minimizes the functional if and only if it is a Gibbs measure.
@@
For any activity $z>0$ and any inverse temperature $\beta>0$, there exists at least one infinite-volume Gibbs measure for the model defined in 2.1.
@@
Moreover, we conjecture that Theorem 3.1 still holds if one makes minor modifications to the model such as
@@ 4.2. Tempered configurations and measures
Let $P\in\mathcal{P}_{\Theta}$ be an infinite-volume Gibbs measure $P\in\mathcal{G}_{z,\beta}(H)$, then $P$ is tempered.
@@ 4.3. Existence of the pressure
$\liminf_{n\to\infty}\frac{\log\big{(}\pi_{\Lambda_{n}}\big{[}\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}+\log(z)N_{\Lambda_{n}}}\big{]}\big{)}}{|\Lambda_{n}|}=\liminf_{n\to\infty}\frac{\log\big{(}\pi_{\Lambda_{n}}\big{[}\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}+\log(z)N_{\Lambda_{n}}}\big{]}\big{)}}{|\Lambda_{n}|}$
and
$\limsup_{n\to\infty}\frac{\log\big{(}\pi^{z}_{\Lambda_{n}}\big{[}\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\big{]}\big{)}}{|\Lambda_{n}|}=\limsup_{n\to\infty}\frac{\log\big{(}\pi^{z}_{\Lambda_{n}}\big{[}\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\big{]}\big{)}}{|\Lambda_{n}|}.$
After this intermediate step, one can then extend the existence of the pressure to boundary conditions sampled from a tempered measure $P$ with finite energy.
@@ 4.4. The variational principle
To show that the set of minimizers coincides with the infinite-volume Gibbs measures, we need to compare measures $P\in\mathcal{P}_{\Theta}$ with finite-volume Gibbs measures. For this note that in finite volumes $\Lambda_{n}$ we have
@@
So a measure $P\in\mathcal{P}_{\Theta}$ is a minimizer, if and only if its relative entropy density (or excess free energy) vanishes. By controlling the effect of different boundary conditions this identity allows us to show that every infinite-volume Gibbs measure is indeed a minimizer.
@@
Let $P$ be an infinite-volume Gibbs measure for activity $z$ and inverse temperature $\beta$, i.e., $P\in\mathcal{G}_{z,\beta}(H)$, then
@@
As a final step, we need to argue that the Gibbs measures are actually the only minimizers. As usual, this is the harder part of the Gibbs variational principle.
@@
then $P$ is an infinite-volume Gibbs measure for activity $z$ and inverse temperature $\beta$, i.e., $P\in\mathcal{G}_{z,\beta}(H)$.
@@ 5. Preliminaries
$H(\omega)=\Big{|}\bigcup_{\boldsymbol{x}\in\omega}B(x,R_{x}+r)\Big{|}\geq\Big{|}\dot{\bigcup_{\boldsymbol{x}\in\omega}}B(x,R_{x})\Big{|}=\sum_{\boldsymbol{x}\in\omega}\left\lvert B(x,R_{x})\right\rvert\gtrsim_{d}\sum_{\boldsymbol{x}\in\omega}R_{x}^{d}.$
Now, for Item (4), we have that
@@
[…]
@@ 6. Energy density
$+\sum_{k\geq 2}\tfrac{1}{k}\sum_{\{\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{k-1}\}\subseteq\omega\setminus(o,R_{x})}\phi_{k}((o,R_{x}),\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{k-1})\mathbbm{1}\Big{\{}x\in\bigcap_{i=1}^{k-1}(\Lambda_{n}-y_{i})\cap\Lambda_{n}\Big{\}}.$
@@
$Q^{o}\big{[}\left\lvert X\right\rvert\mathbbm{1}_{R_{o}>\epsilon^{-1}}\big{]}$ $\leq\epsilon Q^{o}\big{[}\left\lvert X\right\rvert R_{o}\big{]}\lesssim\epsilon Q^{o}\big{[}R_{o}^{d-1}R_{o}\big{]}\lesssim\epsilon.$
@@
where $\Delta^{-}$ is the maximal box inside $\Delta$ such that $\theta_{x}\Delta^{-}\subseteq\Delta$ for all $x\in C$, and we used 5.4. However, by the hardcore condition,
@@
Figure 1. Such a geometric arrangement forces balls with centers outside $\Delta$, which intersect $[-A,A]^{d}\supset B(0,\epsilon^{-1})$, to contain at least one of the $\Delta_{j}$.
@@
i.e., such that we actually only have to look at interactions in the fixed, bounded volume $\Delta$. This can be achieved by choosing, for some scale parameter $A\geq\epsilon^{-1}$, the set $\Delta=[-3A,3A]^{d}$ and decomposing the “boundary” $\Delta\setminus[-A,A]^{d}$ into $M\sim 2^{d}(3^{d}-1)(d-1)^{d/2}$ rectangles with edge length $\sim\frac{A}{2\sqrt{d-1}}$. See Figure 1 for a sketch. This geometric arrangement now has the property that every ball with center outside $\Delta$ which intersects $[-A,A]^{d}\supset B(0,\epsilon^{-1})$ has to fully contain $\Delta_{j}$ for at least one $j\in\{1,\dots,M\}$. The latter property together with the hardcore condition in our model guarantees 6.4. ∎
@@
If infinite energies are allowed, the map $P\mapsto H(P)$ is not continuous. Here is an example of translation-invariant probability measures $(P_{n})_{n\in\mathbb{N}}$ such that $H(P_{n})=\infty$ for all $n\in\mathbb{N}$ and $P_{n}\to P$ in the $\tau_{\mathcal{L}}$-topology, but $H(P)<\infty$. Fix some $S\geq\rho r$ and define the periodic hard-core-respecting configuration
$\omega_{\text{good}}=\sum_{i\in 2S\mathbb{Z}^{d}}\delta_{(i,S)},$
@@
$\omega_{\text{bad},n}=\sum_{i\in 2n\mathbb{Z}^{d}}\delta_{(i,2n)}.$
@@
$\mu_{\text{good}}=S^{-d}\int_{[-S,S]^{d}}\,\delta_{\theta_{y}\omega_{\text{good}}}\,\mathrm{dy}\qquad\text{ and }\qquad\mu_{\text{bad},n}={n}^{-d}\int_{[-n,n]^{d}}\,\delta_{\theta_{y}\omega_{\text{bad},n}}\,\mathrm{dy},$
and set $P_{n}=(1-1/n)\mu_{\text{good}}+(1/n)\mu_{\text{bad},n}$. Then, we have that $H(P_{n})\geq(1/n)H(\mu_{\text{bad},n})=\infty$, while on the other hand $H(\mu_{\text{good}})<\infty$ and $P_{n}\xrightarrow[n\to\infty]{}\mu_{\text{good}}=:P$ in the $\tau_{\mathcal{L}}$-topology.

### v2

@@ 1. Introduction
The Gibbs variational principle is a cornerstone in the equilibrium theory of spatially interacting particle systems. In a nutshell, according to the laws of thermodynamics, equilibrium states of a system (subject to thermodynamic principles) must minimize the excess free energy, which reflects the competition between minimizing energy and maximizing entropy.
@@
On the other hand, equilibrium systems of interacting particles are often described by their conditional finite-volume distributions, in terms of Boltzmann weights with respect to some boundary conditions. Under a consistency condition on these finite-volume measures, one can then hope to find infinite-volume measures that are still consistent with the finite-volume description, and are able to properly describe the macroscopic behavior of the system. This infinite-volume consistency is described via the so-called DLR equations, and ensuring the existence of such measures is one the key tasks in statistical mechanics. In particular, the DLR equations can potentially be solved by multiple infinite-volume measures; this non-uniqueness of equilibrium states constitutes a type of first-order phase transition. However, it is not automatic for a measure to solve the DLR equations and also minimize the excess free energy.
@@
In the case of lattice systems, under the assumption of translation invariance, this link between the DLR equations and the thermodynamic description of equilibria has been established under quite general assumptions for a very large class of models, see [Geo11, Pre76, LR69]. However, moving towards continuum systems, where particles are placed in Euclidean space and even particle numbers in fixed volumes are allowed to fluctuate, the literature becomes much more sparse. In [Geo94, Geo95] the Gibbs variational principle is derived for bounded and unbounded pair potentials in the context of large-deviation results, and with the goal to establish equivalence of ensembles. In particular, no higher-order interactions are allowed, and the particles may not carry individual marks that would for example allow to model interactions between particles of varying sizes. Next, in [DG09] the Gibbs variational principle is established for a particular class of planar models in which (unmarked) particles interact via their associated Delaunay triangulation. More recently, [Der16] considers the case of Gibbs point processes with a finite-range interaction and where the existence of the energy density is guaranteed. However, establishing the existence of the energy density is not always easy and needs to be shown specifically for the model at hand. Moreover, many interesting models do not have a finite-range interaction, see, for example, [Der09, DV20, DV23, KPR12, Lew22, RZ20, Zas22] and the references therein.
@@
The contribution of the present manuscript is therefore to provide a first example of a Gibbs variational principle in the continuum for a system of particles with unbounded marks and with a multibody infinite-range interaction in which also an unbounded number of particles may jointly interact.
@@
We focus on the Asakura–Oosawa model with depletion interaction [AO54, AO58, Vri76], which may be informally described as follows. Imagine a volume with a random number of balls of a random size. The balls cannot penetrate each other, leading to a hardcore repulsion. Additionally, there are second-class particles with a fixed size which may penetrate each other without restriction, but not the balls of the first type. Integrating out the Poisson point process of second-class particles, it can be observed that the balls of type one must additionally obey an area interaction: for them to take up more space comes with the additional cost that, in that space, no second-class particles are present. For more details on the relevance of this colloid-polymer mixture model, see [Vri76, BVS14, RED00, LTV24]. Let us mention that this model can be seen as a generalization of the Widom–Rowlinson model [WR70, CCK95, DH19, DH21, JK17] with an additional internal hardcore. The existence of infinite-volume Gibbs measures (in the sense of the DLR equations) for the Asakura–Oosawa model can be deduced from the criteria presented in [RZ20]. Recently, the model has been analyzed more extensively, for example via cluster-expansion techniques [JT20a, JT20b, JLW22] and with respect to the corresponding Langevin dynamics [FKRZ24, FZ25].
@@ 2.1. Point process formalism
$N_{\Lambda}^{f}:=N_{\Lambda,[0,\infty)}^{f},\quad N_{\Lambda,A}:=N_{\Lambda,A}^{1},\quad\text{ and }\quad N_{\Lambda}:=N_{\Lambda}^{1}.$
@@
where $\mathfrak{B}_{b}$ denotes the set of bounded Borel sets. We also fix some special sets of the position space $\mathbb{R}^{d}$, which will be denoted by
$\Lambda_{n}:=\left[-n/2,n/2\right)^{d},\ n\in\mathbb{N},\quad\text{ and }\quad C:=\Lambda_{1}$
@@ 2.2. Colloidal multi-body interaction
In other words, the interaction models colloidal particles of radius $R_{i}$ centered at $x_{i}$ in a bath of polymers of size $r$. Note that, by inclusion-exclusion, the area-interaction term can be rewritten in terms of $k$-body potentials as
$H^{\rm ar}(\omega)=\sum_{k\geq 1}\,\,\sum_{1\leq i_{1}<\ldots<i_{k}\leq n}\phi_{k}(\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{k}),$
@@ 2.3. Gibbsian formalism
Based on the energy, we now define the finite-volume equilibrium Gibbs measures with respect to suitable boundary conditions. For this, let $\pi^{z}_{\Lambda}$ denote the iid-marked Poisson point process on $\Lambda\times[0,\infty)$ with intensity measure $z\mathds{1}\{x\in\Lambda\}\mathrm{d}x\otimes\mathfrak{R}(\mathrm{d}R)$, with $\Lambda\in\mathfrak{B}(\mathbb{R}^{d})$ and $\mathfrak{R}$ a probability measure on $0,\infty)$. We write $\pi^{z}:=\pi^{z}_{\mathbb{R}^{d}}$. Throughout the paper, we make the following integrability assumption,
$\int\mathrm{e}^{R^{d+\delta}}\mathfrak{R}(\mathrm{d}R)<\infty\qquad\text{ for some }\delta>0.$
@@
Let us note that, since $H_{\Lambda,\zeta}\geq 0$ and $H_{\Lambda,\zeta}(\emptyset)=0$, we have that $\exp(-z|\Lambda|)\leq Z_{\Lambda,z,\beta,\zeta}\leq 1$ and thus the finite-volume Gibbs measures are well-defined for all boundary conditions. One can interpret the random radii $R_{i}$ as annealed disorder in the Gibbs measure.
@@
A probability measure $P\in\mathcal{P}$ is called an infinite-volume Gibbs measure for $z,\beta>0$ if the following DLR equation holds for any bounded $\Lambda\in\mathfrak{B}(\mathbb{R}^{d})$ and any measurable function $f\geq 0$,
@@
We write $\mathcal{G}^{z,\beta}(H)$ for the set of all infinite-volume Gibbs measures for $z,\beta>0$. We denote by $\mathcal{G}^{z,\beta}_{\Theta}(H)=\mathcal{G}^{z,\beta}(H)\cap\mathcal{P}_{\Theta}$ the set of all translation-invariant infinite-volume Gibbs measures.
In the following, the infinite-volume Gibbs measure we consider will always be translation-invariant.
Let us already note that our main result will in particular establish existence of infinite-volume translation-invariant Gibbs measures for the model defined in 2.1. However, the existence could also be verified by checking the conditions presented in [RZ20].
@@ 2.6. The pressure
$\psi(z,\beta):=\lim_{n\to\infty}\frac{\log(\pi^{z}_{\Lambda_{n}}\left[\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\right])}{|\Lambda_{n}|}.$
Here, the existence of the limit $\lim_{n\to\infty}\frac{\log(\pi^{z}_{\Lambda_{n}}\left[\mathrm{e}^{-\beta H_{\Lambda_{n},\zeta}}\right])}{|\Lambda_{n}|}$ is also quite subtle and depends on the choice of the boundary condition $\zeta$. We will show that, for a class of sufficiently nice boundary conditions (including the free case $\zeta=\emptyset$), it exists and actually does not depend on the boundary condition itself.
@@ 3. Main result
    A measure $P\in\mathcal{P}_{\Theta}$ minimizes the functional if and only if it is a translation-invariant Gibbs measure, i.e., $P\in\mathcal{G}^{z,\beta}_{\Theta}$.
@@
For any activity $z>0$ and any inverse temperature $\beta>0$, there exists at least one translation-invariant infinite-volume Gibbs measure for the model defined in 2.1.
@@
Moreover, we conjecture that Theorem 3.1 still holds if one makes minor modifications to the model such as:
@@ 4.2. Tempered configurations and measures
Let $P\in\mathcal{P}_{\Theta}$ be an infinite-volume translation-invariant Gibbs measure $P\in\mathcal{G}_{\Theta}^{z,\beta}(H)$, then $P$ is tempered.
@@ 4.3. Existence of the pressure
$\lim_{n\to\infty}\frac{\log\big(\pi^{z}_{\Lambda_{n}}\big[\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\big]\big)}{|\Lambda_{n}|}=\lim_{n\to\infty}\frac{\log\big(\pi^{z}_{\Lambda_{n}}\big[\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\big]\big)}{|\Lambda_{n}|}.$
After this intermediate step, one can then extend the existence of the pressure to boundary conditions sampled from a tempered measure $P$ with finite energy density.
@@ 4.4. The variational principle
To show that the set of minimizers coincides with the translation-invariants infinite-volume Gibbs measures, we need to compare measures $P\in\mathcal{P}_{\Theta}$ with finite-volume Gibbs measures. For this note that in finite volumes $\Lambda_{n}$ we have
@@
So a measure $P\in\mathcal{P}_{\Theta}$ is a minimizer, if and only if its relative entropy density (or excess free energy) vanishes. By controlling the effect of different boundary conditions this identity allows us to show that every infinite-volume translation-invariant Gibbs measure is indeed a minimizer.
@@
Let $P$ be a translation-invariant infinite-volume Gibbs measure for activity $z$ and inverse temperature $\beta$, i.e., $P\in\mathcal{G}_{\Theta}^{z,\beta}(H)$, then
@@
As a final step, we need to argue that the translation-invariant Gibbs measures are actually the only minimizers. As usual, this is the harder part of the Gibbs variational principle.
@@
then $P$ is a translation-invariant infinite-volume Gibbs measure for activity $z$ and inverse temperature $\beta$, i.e., $P\in\mathcal{G}_{\Theta}^{z,\beta}(H)$.
@@ 5. Preliminaries
$H^{\mathrm{ar}}(\omega)=\Big|\bigcup_{\boldsymbol{x}\in\omega}B(x,R_{x}+r)\Big|\geq\Big|\dot{\bigcup_{\boldsymbol{x}\in\omega}}B(x,R_{x})\Big|=\sum_{\boldsymbol{x}\in\omega}\left\lvert B(x,R_{x})\right\rvert\gtrsim_{d}\sum_{\boldsymbol{x}\in\omega}R_{x}^{d}.$
Now, for Item (4), we have that for $\boldsymbol{x}\in\omega$ it holds
@@
$H^{\mathrm{ar}}(\omega)=\Big|\bigcup_{\boldsymbol{x}\in\omega}B(x,R_{x}+r)\Big|=\int\,\mathrm{d}z\,\mathbbm{1}\{z\in\bigcup_{\boldsymbol{x}\in\omega}B(x,R_{x}+r)\}$
@@ 6. Energy density
$\hskip-28.45274pt+\sum_{k\geq 2}\tfrac{1}{k}\sum_{\{\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{k-1}\}\subseteq\omega\setminus(o,R_{x})}\hskip-22.76228pt\phi_{k}((o,R_{x}),\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{k-1})\mathbbm{1}\Big\{x\in\bigcap_{i=1}^{k-1}(\Lambda_{n}-y_{i})\cap\Lambda_{n}\Big\}.$
@@
$Q^{o}\big[\left\lvert X\right\rvert\mathbbm{1}_{R_{o}>\epsilon^{-1}}\big]$ $\leq\epsilon Q^{o}\big[R_{o}\left\lvert X\right\rvert\mathbbm{1}_{R_{o}>\epsilon^{-1}}\big]\lesssim\epsilon Q^{o}\big[R_{o}R_{o}^{d-1}\big]\lesssim\epsilon.$
@@
where $\Delta^{-}$ is the maximal box inside $\Delta$ such that $\theta_{x}\Delta^{-}\subseteq\Delta$ for all $x\in C=\Lambda_{1}$, and we used 5.4. However, by the hardcore condition,
@@
Figure 1. Such a geometric arrangement forces balls with centers outside $\Delta$, which intersect $A,A]^{d}\supset B(0,\epsilon^{-1})$, to contain at least one of the $\Delta_{j}$.
@@
i.e., such that we actually only have to look at interactions in the fixed, bounded volume $\Delta$. This can be achieved by choosing, for some scale parameter $A\geq\epsilon^{-1}$, the set $\Delta=[-3A,3A]^{d}$ and decomposing the “boundary” $\Delta\setminus[-A,A]^{d}$ into $M\sim 2^{d}(3^{d}-1)(d-1)^{d/2}$ rectangles with edge length $\sim\frac{A}{2\sqrt{d-1}}$. See Figure 1 for a sketch. This geometric arrangement now has the property that every ball with center outside $\Delta$ which intersects $A,A]^{d}\supset B(0,\epsilon^{-1})$ has to fully contain $\Delta_{j}$ for at least one $j\in\{1,\dots,M\}$. The latter property together with the hardcore condition in our model guarantees 6.4. ∎
@@
If infinite energies are allowed, the map $P\mapsto H(P)$ is not continuous. Here is an example of translation-invariant probability measures $(P_{n})_{n\in\mathbb{N}}$ such that $H(P_{n})=\infty$ for all $n\in\mathbb{N}$ and $P_{n}\to P$ in the $\tau_{\mathcal{L}}$-topology, but $H(P)<\infty$. Define the periodic hard-core-respecting configuration
$\omega_{\text{good}}=\sum_{i\in\mathbb{Z}^{d}}\delta_{(i,1/3)},$
@@
$\omega_{\text{bad}}=\sum_{i\in\mathbb{Z}^{d}}\delta_{(i,2)}.$
@@
$\mu_{\text{good}}=2^{-d}\int_{[-1,1]^{d}}\,\delta_{\theta_{y}\omega_{\text{good}}}\,\mathrm{dy}\qquad\text{ and }\qquad\mu_{\text{bad}}={2}^{-d}\int_{[-1,1]^{d}}\,\delta_{\theta_{y}\omega_{\text{bad}}}\,\mathrm{dy},$
and set $P_{n}=(1-1/n)\mu_{\text{good}}+(1/n)\mu_{\text{bad}}$. Then, we have that $H(P_{n})\geq(1/n)H(\mu_{\text{bad}})=\infty$, while on the other hand $H(\mu_{\text{good}})<\infty$ and $P_{n}\xrightarrow[n\to\infty]{}\mu_{\text{good}}=:P$ in the $\tau_{\mathcal{L}}$-topology.